Im not sure how to do this I keep getting a decimal but It shows to be a fraction?
0.15x+0.9(x-1)=0.01(3x-4)
x=?

Answer by fractalier(6550)

Okay, from
0.15x + 0.9(x-1) = 0.01(3x-4)
let's multiply everything by 100 to clear the decimals...
15x + 90(x-1) = 3x - 4
15x + 90x - 90 = 3x - 4
105x - 90 = 3x - 4
102x = 86
x = 86/102 = 43/51